Rene is going to the lake to visit some friends. If the lake is 60 miles away, and Rene is driving at 40 miles per hour the entire time, how long will it take her to get to the lake?
Answer by jim_thompson5910(35256)

d = rt

60 = 40t

60/40 = t

3/2 = t

t = 3/2

t = 1 1/2

So it will take (one and a half) hours. This is equal to 90 minutes.